# Novel inner shadows of the Kerr black hole illuminated by a tilted thin accretion disk

### v2

@@ 4.2 $\Phi=90^{\circ}$ case
Figure 6: Evolution of the inner shadow of a Kerr black hole illuminated by a tilted thin accretion disk with respect to disk inclination $\sigma$ and spin parameter $a$. From left to right, the values of disk inclination are $15^{\circ}$, $30^{\circ}$, $45^{\circ}$, and $60^{\circ}$; from top to bottom, the spin parameters are 00, $0.54$, $0.94$, and $0.9985$. Here, the observation angle and azimuth are fixed at $17^{\circ}$ and $90^{\circ}$, respectively. The field of view and resolution are consistent with the settings in figure 2.
We numerically simulated the inner shadow of a Kerr black hole illuminated by a tilted accretion disk under viewing angles of $17^{\circ}$, $50^{\circ}$, and $85^{\circ}$, with the results presented in figures 6-8, respectively. Similar to the case of $\Phi=0$, we observe that the tilted accretion disk can obscure the inner shadow, leading to a size significantly smaller than $S^{\textrm{min}}_{\Phi=0}$ in certain parameter spaces. Moreover, the introduction of the disk tilt $\sigma$ results in more complex inner shadow structures. Notably, we identify the appearance of dual shadows, as illustrated in figures 7(h), 8(d), and the second row in figure 8. This structure consists of a ‘primary shadow’ located closer to the image center and a ‘secondary shadow’ that is relatively distant but non-negligible. The shape and size of the secondary shadow, often eyebrow-shaped, are influenced by both the spin parameter and the inclination of the accretion disk, with the secondary shadow being more likely to appear for moderate spin values. It is worth mentioning that such dual-shadow structures are unlikely to occur in the equatorial disk scenario. Additionally, while dual-shadow structures can also emerge in binary black hole systems and in the single black hole in vacuum [77, 78, 79, 80], the dual-shadow structure presented in this paper is distinct from those previously reported. Hence, this structure has the potential to serve as a probe for tilted accretion disks.
@@ 5 Conclusions and Discussions
An inner shadow emerges in the black hole image as the inner boundary of the accretion disk intersects the event horizon. We find that in the case of an equatorial accretion disk, the morphology of the inner shadow of a Kerr black hole is shaped by both the viewing angle and the spin parameter, with its size increasing with the former and decreasing with the latter. Thus, when the viewing angle is $0^{\circ}$ and the spin parameter approaches $1$, the Kerr black hole illuminated by the equatorial accretion disk exhibits a minimum inner shadow size of approximately $S_{\textrm{min}}=13.075$ M2.
@@ Acknowledgments
This research has been supported by the National Natural Science Foundation of China [Grant No. 12403081].

### v3

@@ 4.2 $\Phi=90^{\circ}$ case
Figure 9: Evolution of the inner shadow of a Kerr black hole with a tilted thin accretion disk with respect to disk inclination $\sigma$ and spin parameter $a$. From left to right, the values of disk inclination are $15^{\circ}$, $30^{\circ}$, $45^{\circ}$, and $60^{\circ}$; from top to bottom, the spin parameters are 00, $0.54$, $0.94$, and $0.9985$. Here, the observation angle and azimuth are fixed at $17^{\circ}$ and $90^{\circ}$, respectively. The field of view and resolution are consistent with the settings in figure 2.
We numerically simulated the inner shadow of a Kerr black hole with a tilted accretion disk under viewing angles of $17^{\circ}$, $50^{\circ}$, and $85^{\circ}$, with the results presented in figures 9-11, respectively. Similar to the case of $\Phi=0$, we observe that the tilted accretion disk can obscure the inner shadow, leading to a size significantly smaller than $S^{\textrm{min}}_{\Theta=0}$ in certain parameter spaces. Moreover, the introduction of the disk tilt $\sigma$ results in more complex inner shadow structures. Notably, we identify the appearance of dual shadows, as illustrated in figures 10(h), 11(d), and the second row in figure 11. This structure consists of a ‘primary shadow’ located closer to the image center and a ‘secondary shadow’ that is relatively distant but non-negligible. The shape and size of the secondary shadow, often eyebrow-shaped, are influenced by both the spin parameter and the inclination of the accretion disk, with the secondary shadow being more likely to appear for moderate spin values. It is worth mentioning that such dual-shadow structures are unlikely to occur in the equatorial disk scenario. Additionally, while dual-shadow structures can also emerge in binary black hole systems and in the single black hole in vacuum [80, 81, 82, 83], the dual-shadow structure presented in this paper is distinct from those previously reported. Hence, this structure has the potential to serve as a probe for tilted accretion disks.
@@ 5 Conclusions and Discussions
An inner shadow emerges in the black hole image as the inner boundary of the equatorial accretion disk intersects the event horizon. This mechanism remains effective in the tilted accretion disk scenario. In this study, we employ a geometric approach to define a thin, tilted accretion disk and utilize a ray-tracing algorithm to simulate the inner shadow of a Kerr black hole. We find that in the case of an equatorial accretion disk, the morphology of the inner shadow of a Kerr black hole is shaped by both the viewing angle and the spin parameter, with its size increasing with the former and decreasing with the latter. Thus, when the viewing angle is $0^{\circ}$ and the spin parameter approaches $1$, the Kerr black hole surrounded by the equatorial accretion disk exhibits a minimum inner shadow size of approximately $S_{\textrm{min}}=13.075$ M2.
@@
Moreover, it is crucial to emphasize that the tilted accretion disk assumed in this work serves purely as a geometrically defined light source, and the derived inner shadow remains theoretical in nature. From an astrophysical standpoint, the radiation from accreting matter near the black hole’s event horizon is subject to both Doppler redshift and gravitational redshift. This may creates an extremely faint region around the inner shadow that is virtually indistinguishable from the shadow itself, resulting in an observed inner shadow size slightly larger than theoretical predictions. We will investigate this discrepancy in detail in future work.
@@ Acknowledgments
The authors are very grateful to the referee for insightful comments and valuable suggestions. This research has been supported by the National Natural Science Foundation of China [Grant No. 12403081].